OpenStudy (anonymous):

Rational exponent help 4^-3

OpenStudy (anonymous):

1/64

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

Thank you! Could you help me with my other question? :)

OpenStudy (saifoo.khan):

post it.

OpenStudy (anonymous):

Where do I go from here? (x^3)^3/4= x^3/1*^3/4= x^9/4= ^4√x^9

OpenStudy (saifoo.khan):

that is the final answer.

OpenStudy (anonymous):

The question is multiple choice and the options are: a. x√x b. x^2 ^4√x c. x^2 ^4√x^3 d. x^3 ^4√x^3

OpenStudy (saifoo.khan):

d

OpenStudy (saifoo.khan):

that's the same thing, written in different form.

OpenStudy (anonymous):

Hmm, I got it wrong. The correct answer was B. Thanks anyways!

OpenStudy (saifoo.khan):

what? really? was the question u wrote correct?

OpenStudy (anonymous):

\[4^{-3}=\frac{1}{4^{3}}=\frac{1}{164}\]